OpenStudy (anonymous):

can someone please explain to me when you would use an open circle or a closed circle, and when you would shade to the right or left, to graph an inequality on the number line.

OpenStudy (mertsj):

\[< or > means open circle\]

OpenStudy (mertsj):

\[\ge or \le means closed circle\]

OpenStudy (mertsj):

\[> and \ge shade \to the \right\]

OpenStudy (mertsj):